OpenStudy (truwhovian11):

What number is missing from the table of equivalent ratios? A) 2 B) 5 C) 7 D) 13 my guess is B @x__hazel__x

OpenStudy (x__hazel__x):

Are you sure that this is complete?

OpenStudy (truwhovian11):

x y 3 6 ? 10 7 14

OpenStudy (x__hazel__x):

Okay, thank you!

OpenStudy (anonymous):

x = y/2

OpenStudy (anonymous):

So which option when you plug in 10 gives you x?

OpenStudy (milkdud0000):

5 and 7

OpenStudy (x__hazel__x):

Well, in this, you can see that from 3 : 6 it turned to ? : 10, which means that we added 4 to each turn, so add 4 to each one :3

OpenStudy (anonymous):

@x__hazel__x No. @milkdud0000 Not 7.

OpenStudy (milkdud0000):

3 x what equals 6 10 divided by 2 equals 14 divided by 2

OpenStudy (x__hazel__x):

For the left side, you start with 3 and finish with 7, so that means that we added 2

OpenStudy (truwhovian11):

what *cries* i dont understand

OpenStudy (anonymous):

Just do x = y/2 What option is y/2 when y = 10?

OpenStudy (milkdud0000):

you answer is A

OpenStudy (x__hazel__x):

correct.

OpenStudy (truwhovian11):

ok thanks

OpenStudy (truwhovian11):

i ahve more and guess what

OpenStudy (anonymous):

It. Isn't. A.

OpenStudy (milkdud0000):

what is it then

OpenStudy (anonymous):

B. =_=

OpenStudy (milkdud0000):

yes it is B sorry

OpenStudy (x__hazel__x):

oml guys wow. XD

OpenStudy (milkdud0000):

i didnt see the 7

OpenStudy (x__hazel__x):

=.=

OpenStudy (x__hazel__x):

okay..

OpenStudy (truwhovian11):

ok ill medal @Zelda and ill fan evryone sound good

OpenStudy (x__hazel__x):

You already fanned mee <3 ;3

OpenStudy (x__hazel__x):

haha

OpenStudy (milkdud0000):

ok

OpenStudy (truwhovian11):

wait if it was b then my guess was correct @milkdud0000 @x__hazel__x @Zelda

OpenStudy (x__hazel__x):

yes.

OpenStudy (milkdud0000):

yup